raA·2017/18 A: A linear model to predict the Costs of repairs of Toyota station wagon (in S) from its Age in weeks) was fit to 57 cars The model was: Costs-626 +7.3 x Age a) what il the explanatory variable? What do you predict the costs to be for a car aged 100 weeks? A: b) 6. Some 1Q tests are standardized to a Normal model with a mean of 100 and a standard deviation of 16. In what interval would you expect the central 95% of IQ scores to be found? (use the 68-95-99.7 rulc) A: For a sales promotion, the manufacturer places winning symbols under the caps of 30% of all Cola botles If you buy a two-pack of Cola, what is the probability that you win something? 7 A: 8. A broker has calculated the expected values of two different financial instruments X and Y Suppose that EX)-S100, E(Y)-S150. What is the expected value of X Y? A:
Answer #1

4. 1st A: -0.977 , 2nd A: 0.716, 3rd A: -0.021

5. a) The explanatory variable: Age.

b) Costs=-626+7.3* Age=-626+7.3*100=104.

6. Here mean=100 and standard deviation (sd)=16.

The interval is (mean-2*sd,mean+2*sd)=(100-2*16,100+2*16)=(68,32)

7. Here p=0.30 and n=2 . Using Binomial distribution, Probability that you win something is

P(X > 1)-1-P(X < 1-1-P(X = 0) 1ー 0.300(1-0.30)2-0 = 1-0.49 = 0.51

8. E(X)=100, E(Y)=150 then

E(X+Y)=E(X)+E(Y)=100+150=250

E(X-Y)=E(X)-E(Y)=100-150=-50